OpenStudy (anonymous):

(10x^6+20x^4-15x^2) ÷5x^2

OpenStudy (anonymous):

factor out the top part.

OpenStudy (anonymous):

\[\frac{1}{5x^2}10x^6+20x^4-15x^2\]

OpenStudy (anonymous):

or just divide each term separately by dividing the numbers by 5 and subtracting the exponents

OpenStudy (anonymous):

5x^2 cancels at top and bottom leaving 2x^4 + 4x^2+3

OpenStudy (anonymous):

if someone comes up with the answer jus put it on here please!

OpenStudy (anonymous):

for example 10/5 = 2 and \[\frac{x^6}{x^2}=x^{6-2}=x^4\] so first term is \[2x^4\] and work your way down the line

OpenStudy (anonymous):

\[\frac{10x^6}{5x^2}=2x^4\] \[\frac{20x^4}{5x^2}=4x^2\] \[\frac{-15x^2}{x^2}=-3\]